# Inflationary Gravitational Wave Spectral Shapes as test for Low-Scale Leptogenesis

### v1

@@ VI.1 Example spectra for different cases
Fig. 5 illustrates the damped gravitational wave (GW) spectra for Case (a) discussed in Sec. IV, assuming spectral index $n_{T}=0.5$ and $r=0.035$, demonstrating the influence of $T_{\phi}$ and $M_{\phi}$, alongside sensitivity curves of current and future GW experiments. The blue violin lines are taken from recent NANOGrav 15yr dataset NANOGrav:2023hvm ; NANOGrav:2023gor . Different horizontal lines on the plot are coming from the current or projected $\Delta N_{\rm eff}$ bounds from BBN, Planck 2018 and CMB surveys Berbig:2023yyy . In Case (a) scenario, characterized by the condition $T_{\phi}<T_{N_{1}}$, we set the minimum $T_{\phi}$ to be approximately 100 GeV to accommodate sphaleron effects within the electroweak scale. The gray solid line marked “No IMD” is the standard spectrum with no intermediate matter dominated phase while the black solid and dashed lines are the spectral shapes in presence of the matter dominated epoch. Higher frequencies in the figure correspond to earlier times, showcasing suppression beyond the critical frequency $f_{\rm sup}$ due to entropy injection given by Eq. (55). Suppression at frequencies higher than $f_{\rm sup}\sim 10^{-5}-0.1$ Hz are attributed to scalar and right-handed neutrino (RHN) decay, while suppression beyond $\sim 10$ Hz is a consequence of inflaton decay terminating at the reheating temperature $T_{\rm RH}=10^{10}$ GeV. Lower values of $T_{\phi}$ shift the suppression frequency to the left and increase dilution factor. The solid and dashed spectral lines represent $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively. A higher $M_{\phi}$ means more dilution. Importantly, GW signatures remain invariant across weak and strong washout regimes, as they solely depend on the scalar dynamics in this case. Fig. 6 is same as Fig. 5 but shown for $n_{T}=0$.
Figure 12: Same as Fig. 10 but for $T_{\phi}=100$ GeV and $n_{T}=0.5$, $M_{\phi}=10^{15}$ GeV. Here $m_{*}=1.1\times 10^{-3}$ eV.
Fig. 7 presents the GW signature of long-lived RHNs in Case (b) for different values of the washout parameter $K$ and lightest RHN mass $M_{1}$. Once again the solid and dashed spectral lines correspond to $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively. We choose the combination $T_{N_{1}}=M_{1}\sqrt{K}$ in a way such that $100\;\text{GeV}\leq T_{N_{1}}<T_{\phi}$, where we have taken $T_{\phi}=10^{7}$ GeV. Other parameters are set as: $r=0.035,n_{T}=0.5,T_{\rm RH}=10^{10}$ GeV. We see that the role of $T_{\phi}$ seen in Fig. 5 is taken up by $T_{N_{1}}$ in this case, as expected.
Case (c) mirrors thermal leptogenesis along with 2-step entropy injection as discussed earlier. Figures 8 and 9 depict the GW signatures of this scenario. For these plots, we make sure that the thermalization condition given in Eq. (84) is satisfied assuming $y_{N_{i}}=y_{R}$, such the two-step entropy injection scenario is applicable. We use the 2-step transfer function utilizing Eq. (57). Notably, the scalar’s effect manifests as an additional dip in the spectrum, a knee-like feature, positioned between the dips induced by inflaton decay and RHN decay. Thus we have two knee-like features, one from $\phi$ decay and another from RHN decay. In Fig. 8 we show the dependency of the spectral shape on $M_{1}$ and $K$, fixing $T_{\phi}=10^{5}$ GeV, $M_{\phi}=10^{13}$ GeV, $T_{\rm RH}=10^{15}$ GeV for $n_{T}=0.5$. The solid and dashed lines correspond to $M_{1}=10^{5}$ and $3\times 10^{5}$ GeV respectively. Note that for $T_{N_{1}}<100$ GeV, efficient leptogenesis is not possible as sphalerons are turned off.
Finally in Fig. 9, we show the dependency of the spectrum on $T_{\phi}$ and $M_{\phi}$ fixing $M_{1}=10^{6}$ GeV, $K=10^{-8}$ such that leptogenesis scale is $T_{N_{1}}=100$ GeV. Other parameters are same as in Fig. 8. The solid and dashed lines here represent $M_{\phi}=10^{13}$ and $3\times 10^{13}$ GeV respectively. The entropy injection from scalar is independent of leptogenesis scale but it reduces the projected SNR for different experiments via the additional suppression as seen in the plots.
@@
Figure 14: Same as Fig. 10 but demonstrating thermal leptogenesis (Case (c) in Sec. IV) with two-step entropy injection for $n_{T}=0$ (Left) and $n_{T}=0.5$ (Right). Here $m_{*}=1.1\times 10^{-3}$ eV. For both plots we took $M_{\phi}$= $10^{15}$ GeV , $T_{\phi}$ = $10^{10}$ GeV, $r=0.035$. The green shaded region represents instantaneous RHN decay (Case (a)). Two-step entropy injection is not possible in the brown shaded region as $T_{\phi}$ is less than RHN domination temperature $T_{\rm dom}^{N_{1}}\sim 2\%M_{1}$. See main text for detailed description and analysis.
Figure 15: Same as Fig. 14 but for $T_{\phi}=10^{6}$ GeV.
VI.2 Estimates for Signal to Noise Ratio
The next set of Figures 1011, 12131415 demonstrates the projected future sensitivities of U-DECIGO, DECIGO, BBO, $\mu$-ARES, LISA, CE and ET experiments in terms of signal to noise ratio (SNR) given in Eq. (66), in the parameter space of the RHN for different values of $n_{T},M_{\phi},T_{\phi},T_{\rm RH}$. Figures 10111213 correspond to non-thermal leptogenesis with black dashed line separating the Case (a) and Case (b) regions. Figures 14 and 15 demostrate the thermal leptogenesis with two-step entropy injection scenario, Case (c), with green region specifying Case (a) and brown region corresponding to $T_{\phi}<T_{\rm dom}^{N_{1}}$, where two-step entropy injection is not possible as described in Sec IV.
The gray shaded region is excluded from the requirement that $T_{\rm eq.}$ should be larger than the BBN temperature of a few MeV. In Fig. 14 and 15 we are not interested in $K\gtrsim 10^{-4}$ where RHN domination is not possible, i.e. Eq. (32) is not satisfied, thus GW is not connected to leptogenesis in this region. In all the SNR plots, the colored regions associated with different experiments can be probed by that experiment with SNR $\geq 10$. In all these SNR plots we choose $T_{\rm RH}=10^{16}$ GeV in order to eliminate its effect on the SNR. In vanilla leptogenesis scenario, vacum stability of SM Higgs gives an upper bound on the lightest RHN mass $M_{1}\lesssim 10^{14}$ GeV Casas:1999cd ; Elias-Miro:2011sqh and we assume it would be approximately the same in our model Fu:2023nrn . Hence we take upper limit on $M_{1}$ to be around $10^{14}$ GeV. On the other hand, the perturbativity limit of the model requires the Yukawa couplings $\lambda_{ij}<4\pi$, which along with the necessity to produce at least one mass eigenstate $m_{\nu}=0.05$ eV gives another bound, $M_{2}\lesssim 10^{15}$ GeV. Hence, in our quasi-degenerate RHN model, we find the perturbativity bound $M_{1}\lesssim 10^{15}$ GeV. The perturbativity limit of the light-heavy neutrino mixing $y_{N_{i}}<4\pi$ is always satisfied in our model.
@@ VI.2 Estimates for Signal to Noise Ratio
In Fig. 10, we pick a high scalar mass and decay temperature $M_{\phi}=10^{15}$ GeV, $T_{\phi}=10^{10}$ GeV and show the projected sensitivities of different experiments for $n_{T}=0$ (Left) and $n_{T}=0.5$ (Right). We see that increasing the spectral index $n_{T}$ also increases the GW sensitive region by increasing the SNR as expected. For $n_{T}=0$, LISA and ET will not be able to probe leptogenesis but for $n_{T}=0.5$ both the experiments will be sensitive to a large portion of the parameter space. In the Case (a) region of instantaneous RHN decay, the SNR will not depend on $M_{1}$ and $K$ but only on $M_{\phi}$ and $T_{\phi}$. In the Case (b) region, RHNs decay at $T_{N_{1}}=M_{1}\sqrt{K}$. Hence decreasing $M_{1}$ or $K$ results in more suppression and lower SNR.
@@
In Fig. 11, we show the sensitivities for $n_{T}=0$, $T_{\phi}$ = $10^{6}$ GeV and $M_{\phi}$= $10^{15}$ GeV (Left) and $10^{10}$ GeV (Right). Both the left and right plots show almost the same sensitive area in the parameter space but suppression in the left plot is expected to be higher due to higher mass scalar decay. In both the plots we see the Case (a) region covering more space due to lower $T_{\phi}$ value compared to Fig. 10. We also notice that CE is no longer sensitive for $n_{T}=0,T_{\phi}=10^{6}$ GeV in Fig. 11 (Left) while it is sensitive for $T_{\phi}=10^{10}$ GeV in Fig. 10 (Left).
In Fig. 12, we show the effect of a low scalar decay temperature $T_{\phi}=100$ GeV assuming $M_{\phi}=10^{15}$ GeV. We note that lowering $T_{\phi}$ increases the area of the region of instantaneous RHN decay (Case (a)) in the parameter space, and for $n_{T}=0.5$ most of the parameter space is accessible to DECIGO, BBO, $\mu-$ARES and U-DECIGO. We find no experiment to be sensitive to leptogenesis for $n_{T}=0$ in this case. Since $T_{\phi}=100$ GeV, below this scale efficient leptogenesis is not possible in the Case (b) region in this figure.
@@
Fig. 13 shows the shape of the sensitive regions for $M_{\phi}=10^{5}$ GeV and $T_{\phi}=100$ GeV assuming $n_{T}=0$. We see that even for $n_{T}=0$, UDECIGO, DECIGO, BBO, $\mu-$ARES can probe leptogenesis with $T_{\phi}=100$ GeV, given small mass $M_{\phi}$ of the scalar. A small scalar mass injects less amount of entropy when they decay as a result of which the suppression is small, hence SNR is larger.
In Fig. 14 and 15 we show the SNR contours in the Case (c) region of two-step entropy injection assuming a non-vanishing coupling $y_{R}$. We make sure that the thermalization condition (84) is satisfied for given parameter choices as well as the condition for two-stages of entropy injection, $T_{\phi}>T_{\rm dom}^{N_{1}}$ described in Sec. IV. In Fig. 14 we chose $M_{\phi}=10^{15}$ GeV, $T_{\phi}=10^{10}$ GeV and $T_{\rm RH}=10^{16}$ GeV and demosntrate the effect of the spectral index $n_{T}=0$ (Left) and $n_{T}=0.5$ (Right). Note that instantaneous RHN decay scenario (Case (a)) is still possible in thermal scenario, denoted by the green shaded region. Fig. 15 is the same as Fig. 14 but for $T_{\phi}=10^{6}$ GeV.
@@
We calculate the SNR contours for all these GW experiments taking various values of $M_{\phi}$ and $T_{\phi}$ for $n_{T}=0$ and $0.5$, which we do not show in this paper. However, based on these results and Fig. 10-15, in Tables 2 and 3, we summarize which key GW experiments can probe the parameter space $M_{1}$ vs $K$, associated to thermal and non-thermal leptogenesis in the future with threshold SNR $\geq 10$, for these parameter choices. We see that thermal leptogenesis is possible only for high scalar mass $M_{\phi}\sim 10^{15}$ GeV which U-DECIGO, BBO and $\mu-$ARES will be able to probe even for $n_{T}=0$ while LISA and ET requires $n_{T}\gtrsim 0.5$.
@@
$10^{15}$ $10^{10}$ Non-Th, Th Non-Th, Th Non-Th, Th - - Non-Th, Th
$10^{6}$ Non-Th, Th Non-Th, Th Non-Th, Th - - -
$10^{2}$
$10^{10}$ $10^{6}$ Non-Th Non-Th Non-Th - - -
$10^{2}$
$10^{5}$ $10^{2}$ Non-Th Non-Th Non-Th - - -
Table 2: Future sensitivity of different key experiments to probe the RHN parameter space ($M_{1}$ vs. $K$) with threshold SNR $>10$, given benchmark values of $M_{\phi}$ and $T_{\phi}$ for $n_{T}=0$. Non-Th and Th implies sensitivity of the specific experiment to non-thermal and thermal leptogenesis. Note that thermal scenario requires Eq. (84) to be satisfied. Here we considered $T_{\rm RH}=10^{16}$ GeV and maximal tensor-to-scalar ratio $r=0.035$.
$M_{\phi}\,$ (GeV) $T_{\phi}\,$ (GeV) U-DECIGO BBO $\mu-$ARES LISA ET CE
$10^{15}$ $10^{10}$ Non-Th, Th Non-Th, Th Non-Th, Th Non-Th, Th Non-Th, Th Non-Th, Th
$10^{6}$ Non-Th, Th Non-Th, Th Non-Th, Th Non-Th, Th Non-Th, Th Non-Th, Th
$10^{2}$ Non-Th Non-Th Non-Th - - -
$10^{10}$ $10^{6}$ Non-Th Non-Th Non-Th Non-Th Non-Th Non-Th
$10^{2}$ Non-Th Non-Th Non-Th Non-Th Non-Th Non-Th
$10^{5}$ $10^{2}$ Non-Th Non-Th Non-Th Non-Th Non-Th Non-Th
Table 3: Same as Table. 2 but for $n_{T}=0.5$. We see that most of the parameter space $M_{\phi}$ vs $T_{\phi}$ will be probed by future GW experiments if $n_{T}=0.5$.
VI.3 Compatibility with NANOGrav
Several papers NANOGrav:2023hvm ; Vagnozzi:2023lwo ; Datta:2023xpr ; Ellis:2023oxs discuss the possibility of inflationary gravitational waves to be the source of the stochastic background detected in the recent NANOGrav data NANOGrav:2023gor . The NANOGrav collaboration NANOGrav:2023hvm states the best fit is found to be around $n_{T}\sim 2.61\pm 0.85$, $r\sim 10^{-14.06\pm 5.82}$ and $T_{\rm RH}\sim 1$ GeV. However at such low reheating temperatures, efficient leptogenesis as we discussed in the paper is not possible due to sphaleron constraints.

### v2

@@ VI.1 Example spectra for different cases
Fig. 8 illustrates the damped gravitational wave (GW) spectra (black lines) for Case (a) and Case (b) discussed in Sec. IV, assuming $r=0.035,M_{1}=10^{5}$ GeV, $K=10^{3}$ and spectral index $n_{T}=0$ (Left) and $n_{T}=0.3$ (Right). Here we demonstrate the influence of $T_{\phi}$ and $M_{\phi}$ on the GW spectra. We set the minimum $T_{\phi}$ to be approximately 100 GeV to accommodate sphaleron effects within the electroweak scale. The gray solid line marked “No IMD” is the standard spectrum with no intermediate matter-dominated phase while the black solid and dashed lines are the spectral shapes in the presence of the matter-dominated epoch. Different colored regions correspond to sensitivity curves of current and future GW experiments. The blue violin lines are taken from the recent NANOGrav 15yr dataset NANOGrav:2023hvm ; NANOGrav:2023gor . Various horizontal lines on the top of the plot are coming from the current or projected $\Delta N_{\rm eff}$ bounds from BBN, Planck 2018, and CMB surveys Berbig:2023yyy . Higher frequencies in the figure correspond to earlier times, showcasing suppression beyond the critical frequency $f_{\rm sup}$ due to entropy injection given by Eq. (68). Suppression at frequencies higher than $f_{\rm sup}\sim 10^{-5}-0.1$ Hz are attributed to majoron and RHN decay, while suppression at frequencies higher than $\sim 10$ Hz is a consequence of inflaton decay terminating at the reheating temperature $T_{\rm RH}=10^{13}$ GeV. Lower values of $T_{\phi}$ shift the suppression frequency to the left and increase the dilution factor as expected from Eq. (50). The solid and dashed spectral lines represent $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively. A higher $M_{\phi}$ means more dilution. Importantly, GW signatures remain invariant across weak and strong washout regimes, as they solely depend on the majoron dynamics in this case.
Figure 10: Projected sensitivities of different experiments in the RHN parameter space for $n_{T}=0$ (Left) and $n_{T}=0.3$ (Right). Here $\tilde{m}_{1}$ is the effective neutrino mass defined by Eq. (8) and $m_{*}=1.1\times 10^{-3}$ eV is the equilibrium neutrino mass defined in Eq. (9). For both plots we took $M_{\phi}$= $10^{15}$ GeV , $T_{\phi}$ = $10^{10}$ GeV, $r=0.035$. The colored regions associated with each experiment imply sensitivity with SNR >10>10. As expected, a higher $n_{T}$ increases overall SNR. Non-thermal leptogenesis is possible in this model only in the region between the two solid red lines. See the main text for a detailed description and analysis.
Fig. 9 presents the GW signature of long-lived RHNs in Case (c), depicting the two stages of entropy injection. We show the dependence of the spectra on parameters $K$ and $M_{1}$ (Left) and on $M_{\phi}$ and $T_{\phi}$ (Right). The necessary condition for this case is $\sqrt{K}<2T_{\phi}/M_{\phi}$. We choose the combination $T_{N_{1}}=M_{1}\sqrt{K}$ in a way such that $T_{N_{1}}\geq 100\;\text{GeV}$. For these spectra, we take $r=0.035$, $n_{T}=0.3$, and use the 2-step transfer function utilizing Eq. (70). Notably, the majoron’s effect manifests as an additional dip in the spectrum, a knee-like feature, positioned between the dips induced by inflaton decay and RHN decay. Thus we have two knee-like features, one from $\phi$ decay and another from RHN decay. In the Left plot, the solid and dashed spectral lines correspond to $M_{1}=10^{8}$ and $3\times 10^{8}$ GeV respectively. Here we take $M_{\phi}=10^{12}$ GeV, $T_{\phi}=10^{8}$ GeV and $T_{\rm RH}=10^{13}$ GeV. We see that a lower $K$ and $M_{1}$ lead to higher suppression, as expected from Eq. (52). In the Right plot, we fix $M_{1}=10^{8}$ GeV, $K=10^{-12}$, $T_{\rm RH}=10^{13}$ GeV. Here solid and dashed spectral lines correspond to $M_{\phi}=10^{12}$ and $2\times 10^{12}$ GeV respectively. Here, we see that the overall suppression increases with increasing $T_{\phi}$ and decreasing $M_{\phi}$, also understood from Eq. (52). This trend is opposite to the trend in Cases (a), (b) as shown in Fig. 8.
In Case (c), the suppression factor associated with the the middle “knee-like” feature is given by, $R^{\rm 1st}_{\rm sup}=(\Delta_{\phi}\Delta_{N})^{-4/3}$, defined in Eq. (77). Minimum possible value of this factor corresponds to the maximum suppression. In this case, $R^{\rm 1st}_{\rm sup}\sim 693.5\,K^{2/3}$, obtained by substituting Eq. (50) and (52) in Eq. (77). Minimum value of the washout parameter $K$ is obtained by considering RHN decay temperature $T_{N_{1}}=M_{1}\sqrt{K}\sim 100$ GeV such that sphalerons can successfully convert lepton asymmetry to baryon asymmetry. For maximum value of $M_{1}=10^{14}$ GeV, we obtain the minimum value of $R^{\rm 1st}_{\rm sup}\simeq 7\times 10^{-14}$. If we do not consider leptogenesis, then $T_{N_{1}}$ can go as low as BBN temperature which is around 10 MeV. In that case the minimum suppression factor is $R^{\rm 1st}_{\rm sup}\simeq 3\times 10^{-19}$. If in the future, two-step entropy injection with a smaller suppression factor is detected, it would rule out non-thermal leptogenesis scenario within our model.
VI.2 Estimates for Signal to Noise Ratio
Figure 12: Same as Fig. 11 (Right) but for $n_{T}=0.3$. Here $m_{*}=1.1\times 10^{-3}$ eV.
The next set of Figures 1013 demonstrates the projected future sensitivities of U-DECIGO, DECIGO, BBO, $\mu$-ARES, LISA, CE and ET experiments in terms of signal to noise ratio (SNR) given in Eq. (79), in the parameter space of the RHN for different values of $n_{T},M_{\phi},T_{\phi},T_{\rm RH}$. In these figures, the black dashed lines separate the Case (a), Case (b) and Case (c) regions. In each of the SNR plots, the gray-shaded region is excluded due to the requirement that the latest radiation-dominated era should start before BBN takes place at around temperatures of a few MeV. In our model, the latest radiation domination starts at the equilibrium temperature $T_{\rm eq.}=T_{\rm dec.}$ defined in Eq. (48). In all the SNR plots, the colored regions associated with different experiments can be probed by that experiment with SNR $\geq 10$. In all these SNR plots we choose $T_{\rm RH}=10^{16}$ GeV to eliminate its effect on the SNR. The two solid red lines represent theoretical constraints on our model. In the region above the upper red line, $v_{B-L}$ becomes larger than Planck scale while below the lower red line, the requirement that the confinement scale should be less than the BLB-L breaking scale, $\Lambda<v_{B-L}$, discussed in Sec. III, is not satisfied. Hence, successful leptogenesis is possible only in the region between these two lines in our model. In vanilla leptogenesis scenario, vacuum stability of SM Higgs gives an upper bound on the lightest RHN mass $M_{1}\lesssim 10^{14}$ GeV Casas:1999cd ; Elias-Miro:2011sqh and we assume it would be approximately the same in our model Fu:2023nrn . Hence we take the upper limit on $M_{1}$ to be around $10^{14}$ GeV. On the other hand, the perturbativity limit of the model requires the Yukawa couplings $\lambda_{ij}<4\pi$, which along with the necessity to produce at least one mass eigenstate $m_{\nu}=0.05$ eV gives another bound, $M_{2}\lesssim 10^{15}$ GeV. Hence, in our quasi-degenerate RHN model, we find the perturbativity bound $M_{1}\lesssim 10^{15}$ GeV. The perturbativity limit of the light-heavy neutrino mixing $y_{N_{i}}<4\pi$ is always satisfied in our model.
@@ VI.2 Estimates for Signal to Noise Ratio
In Fig. 10, we pick a high majoron mass $M_{\phi}=10^{15}$ GeV, and decay temperature $T_{\phi}=10^{10}$ GeV and show the projected sensitivities of different experiments for $n_{T}=0$ (Left) and $n_{T}=0.3$ (Right). We see that increasing the spectral index $n_{T}$ also increases the GW-sensitive region by increasing the SNR as expected. For $n_{T}=0$, LISA and ET will not be able to probe leptogenesis but for $n_{T}=0.3$ both the experiments will be sensitive to a large portion of the parameter space. In the Case (a) and Case (b) regions, the SNR will not depend on $M_{1}$ and $K$ but only on $M_{\phi}$ and $T_{\phi}$. In the Case (c) region, RHNs decay at $T_{N_{1}}=M_{1}\sqrt{K}$. Hence decreasing $M_{1}$ or $K$ results in more suppression and lower SNR.
@@
In Fig. 11, we show the sensitivities for $n_{T}=0$, $T_{\phi}$ = $10^{6}$ GeV and $M_{\phi}$= $10^{15}$ GeV (Left) and $10^{10}$ GeV (Right). We see that for $M_{\phi}=10^{10}$ GeV, the GW sensitive region in the parameter space is smaller compared to that for $M_{\phi}=10^{15}$ GeV. Also, comparing Fig. 11 (Left) with Fig. 10, we see that the Case (a) and Case (b) regions cover more space for lower $T_{\phi}$ values. We notice that CE is no longer sensitive for $n_{T}=0,T_{\phi}=10^{6}$ GeV in Fig. 11 (Left) while it is sensitive for $T_{\phi}=10^{10}$ GeV in Fig. 10 (Left).
In Fig. 12, we show the SNR for $n_{T}=0.3$, $T_{\phi}=10^{6}$ GeV and $M_{\phi}=10^{10}$ GeV. Comparing with Fig. 11 (Right) plot, again the effect of $n_{T}$ is seen, and LISA, CE, and ET will be able to probe the parameter space for $n_{T}=0.3$.
@@
Fig. 13 shows the shape of the sensitive regions for $M_{\phi}=10^{5}$ GeV and $T_{\phi}=100$ GeV assuming $n_{T}=0$. We see that UDECIGO, DECIGO, BBO, $\mu$-ARES can probe leptogenesis with $T_{\phi}=100$ GeV, given small mass $M_{\phi}$ of the majoron. A small majoron mass injects less amount of entropy when they decay as a result of which the suppression is small, hence SNR is larger.
@@
We calculate the SNR contours for all these GW experiments taking various values of $M_{\phi}$ and $T_{\phi}$ for $n_{T}=0$ and $0.3$ but do not show all of them in this paper. However, based on these results and Fig. 10-13, we summarize which key GW experiments can probe the parameter space $M_{1}$ vs $K$ in Tables 2.
@@
$n_{T}=0$ $n_{T}=0.3$ $n_{T}=0$ $n_{T}=0.3$ $n_{T}=0$ $n_{T}=0.3$ $n_{T}=0$ $n_{T}=0.3$ $n_{T}=0$ $n_{T}=0.3$ $n_{T}=0$ $n_{T}=0.3$
$^{15} 1010$$^{10}$ - -
$10^{6}$ - - -
$10^{2}$ - - - - - - - - - - -
$10^{10}$ $10^{6}$ - - -
$10^{2}$ - - - - - - - - -
$10^{5}$ $10^{2}$ - - -
Table 2: Future sensitivity of different key GW experiments to probe the RHN parameter space ($M_{1}$ vs. $K$) with threshold SNR >10>10, for benchmark values of $M_{\phi}$ and $T_{\phi}$. Columns show sensitivities for $n_{T}=0$ and $n_{T}=0.3$. Here the check mark implies that the particular experiment will be sensitive to non-thermal leptogenesis for given value of $n_{T}$. Dashes imply that in the cosidered parameter space, SNR <10<10 for the given experiment. We considered $T_{\rm RH}=10^{16}$ GeV and maximal tensor-to-scalar ratio $r=0.035$.
It is worth mentioning that there are certain models of inflation that go beyond the standard power-law description of the IGW spectrum. For example, a broken power-law spectrum of IGW is suggested in Jiang:2023gfe . The power-law is broken at a scale $k_{\rm break}>k_{*}$ corresponding to the frequency $f_{\rm break}\ll 10^{-8}$ Hz, where $k_{*}$ is the pivot scale defined in Eq. (55). Below $f_{\rm break}$, the spectrum is red-tilted ($n_{T}^{\rm CMB}\lesssim 0$), while above $f_{\rm break}$ it can be blue-tilted ($n_{T}>0$). Since our focus is on $f\gg 10^{-8}$ Hz, we can replace $k_{*}$ with $k_{\rm break}$ in Eq.(55) to obtain broken power law assuming $n_{T}^{\rm CMB}=0$. In this case, $\Omega_{\rm GW}(k)$ and the resulting SNR are suppressed by a factor $(k_{*}/k_{\rm break})^{n_{T}}\sim\mathcal{O}(0.1$$1)$ relative to the pure power-law case. Similarly, inflationary models with a Gauss-Bonnet term can yield a blue-tilted spectrum with a running $n_{T}$ Koh:2018qcy , also modifying the SNR. However, the effect of an intermediate matter-dominated era quantified by the ratio of SNR with and without the matter-domination, remains unaffected. We therefore do not consider such extensions further in this work.